OpenStudy (anonymous):

Calculate the unknown side x for each triangle below. 3^2+x^2=6^2 a. b.

OpenStudy (anonymous):

\[b^2=c^2-a^2\]

OpenStudy (anonymous):

just substitute b with x

OpenStudy (anonymous):

then look for the value of b

OpenStudy (anonymous):

@panlac01 let me draw the pic for U

OpenStudy (anonymous):

are we talking about a right triangle here?

OpenStudy (anonymous):



OpenStudy (anonymous):

okay, use the pythagorean theorem

OpenStudy (anonymous):

\[a^2+b^2=c^2\]

OpenStudy (anonymous):

let 6 be the a, 3 be the b and x be the c

OpenStudy (dumbcow):

are you sure? according to the equation, the hypotenuse of triangle should be 6

OpenStudy (anonymous):

@dumbcow and @panlac01 so is it 3^2+x^2=6^2 or 6^2+3^2=x^2

OpenStudy (anonymous):

the latter

OpenStudy (anonymous):

if we followed your drawing then it is the latter

OpenStudy (anonymous):

According to your drawing: \[\color{red}{x^2 = 6^2 + 3^2}\]

OpenStudy (anonymous):

@panlac01 ok so I got 45-x^2=0